OpenStudy (anonymous):

Which graph shows the solution of the inequality? Please Help I'll Give Medal!!!

OpenStudy (anonymous):



OpenStudy (anonymous):

Take your time

OpenStudy (ranga):

\[-\frac{x}{2} \le 8\]To solve for x multiply both sides by 2:\[-x \le 16\]

OpenStudy (anonymous):

ok yeah thats what i got but i don't know how to do the thing on the number line to get my answer

OpenStudy (anonymous):

But i understand that part

OpenStudy (ranga):

Multiply both sides by -1. When you multiply by a negative sign you have to flip the "less than or equal to" to "greater than or equal to". \[ x \ge -16\]

OpenStudy (anonymous):

i thought i multiply by 2 ?

OpenStudy (ranga):

So on a number line, mark the -16 and anything to the right of it including -16 is a solution.

OpenStudy (anonymous):

Not 1?

OpenStudy (ranga):

I multiplied by 2 and got -x <= 16 but I want to solve for x and not minus x. So multiply both sides by -1. x >= -16 is the solution.

OpenStudy (ranga):

When you multiply by a negative number the inequality will flip.

OpenStudy (anonymous):

so your still going to get the same answer?

OpenStudy (ranga):

same as what? The steps I told you is the way to get the correct answer.

OpenStudy (anonymous):

Ohhh the sign will flip the > and <

OpenStudy (anonymous):

Correct?

OpenStudy (ranga):

Yes, multiplying by a NEGATIVE number will flip > to < or < to > or <= to >= or >= to <=

OpenStudy (ranga):

Here we have <= and so multiplying or dividing by a negative number will flip the inequality to >=

OpenStudy (anonymous):

ok

OpenStudy (ranga):

Oh, it lost my drawing. :(

OpenStudy (ranga):

But the solution is x in the interval [-16, infinity) In a number line mark the point -16 and draw a horizontal line from -16 to the positive x side and put an arrow. Make sure that point -16 is included in the solution because we have x >= -16.